OpenStudy (anonymous):

Relationship B has a greater rate than Relationship A. The graph represents Relationship A. Which equations could represent Relationship B? Choose all answers that are correct.

OpenStudy (anonymous):

A. y = 3/4x B. y = 0.6x C. y = 2/3x D. y = 1/4x

OpenStudy (anonymous):

@AloneS

alones (alones):

oh my gosh it's something familiar :/

OpenStudy (anonymous):

but do you get it

alones (alones):

i'm pretty sure you have to use this formula y=mx+b sorry dude i suck at explaing things :/

OpenStudy (anonymous):

do you know anyone who is good at math

alones (alones):

well i do @AlexandervonHumboldt2 his good but ain't suere if his online right now :/

alones (alones):

oh ask @ganeshie8